find the area bounded by the curves y=x|x|, x-axis,x=1 and x=-1.

Grade:12

2 Answers

Sher Mohammad IIT Delhi
10 years ago
y=x^2  for x>0
y=-x^2 for x<0


if we are just checking the magnitude of area,


Sher Mohammad,
B.tech, IIT Delhi

sanjeev100
10 years ago
=((X^3)/3)+((X^3)/3) =(1/3+1/3) =2/3